OpenStudy (anonymous):

What happens to the area of a rectangle if we mul-tiply both its length and width by 2? by 3? by 10? by n?

OpenStudy (hoblos):

original area= A= l*b if we multiply both its length and width by 2 new area = 2l*2b = 4l*b = 4A if we multiply both its length and width by 3 new area = 3l*3b = 9l*b = 9A if we multiply both its length and width by n new area = nl*nb = (n^2)l*b = (n^2)A

OpenStudy (anonymous):

I agree with Hoblos.

Directrix (directrix):

by 2 ----> multiplication by 2^2 or 4 by 3 -->multiplied by 3^2 or 9 by 10 ---> multiplied by 10^2 by n ----> multiplied by n^2